OpenStudy (anonymous):

what is the sum of the terms of the series (-1)+(-2)+(-3)+...+(-10) help !!

OpenStudy (anonymous):

well with just those four numbers it would be -16

OpenStudy (studygurl14):

-55

OpenStudy (studygurl14):

-1 + -2 + -3 + -4 + -5 + -6 + -7 + -8 + -9 + -10 = -55

OpenStudy (anonymous):

oh i see so he was just posting it the faster way... @StudyGurl14 is right

OpenStudy (anonymous):

its arithmetic series, and its actually 10 numbers.

OpenStudy (anonymous):

give her a medal!

OpenStudy (studygurl14):

Also, you can use the following formula...

OpenStudy (anonymous):

can you explain to me how you got -55?

OpenStudy (studygurl14):

where n = number of terms, a1 = first term, a2 = nth term, Sn= sum \(\Large S_n=\frac{n(a_1+a_n)}{2}\)

OpenStudy (studygurl14):

\(\Large S_{10}=\frac{10(-1+-10)}{2}\rightarrow \frac{10(-11)}{2}\rightarrow\frac{-110}{2}=-55\)

geerky42 (geerky42):

Or you can factor -1 out, and you have \(-1(1+2+3+\cdots+10)\) To find the sum fron 1 to n: you have \(S = \dfrac{n(n+1)}{2}\) So you have \(-\dfrac{10(11)}{2} = -\dfrac{110}{2}=\boxed{-55}\)

OpenStudy (studygurl14):

True @geerky42

OpenStudy (anonymous):

thanks for y'all help , i appreciate this a lot. i wish you guys can help me out a lot more with this .

OpenStudy (studygurl14):

:)

OpenStudy (anonymous):

expand. \[\sum_{n=2}^{6}3n\]

geerky42 (geerky42):

Just use formula @StudyGurl14 gave you. \(S_n = \dfrac{n(a_1+a_n)}{2}\) For first term, n=2 so you have \(a_1=3(2)=6\) For final term, n=6 so you have \(a_n = 3(6)=18\) And there are \(6-2+1=5\)terms. So you have \(\dfrac{5(6+18)}{2}=\boxed{60}\)

OpenStudy (anonymous):

where did you get 6-2+1= 5 terms from? @geerky42

OpenStudy (anonymous):

it says expand, its not asking for the sum.

geerky42 (geerky42):

Shortcut to count terms, you just subtract upper limit by lower limit, then add one. Instead of count how many terms 2, 3, 4, 5, 6, you just do math: \(6-2+1=5\) So we know that there are five terms. Useful if you have a ton of terms to count.

OpenStudy (anonymous):

why are you adding 1 , where does that come from?

geerky42 (geerky42):

Well, expanding should be easy. \[\sum_{n=2}^63n = 3(2)+3(3)+3(4)+3(5)+3(6)\]I am sure you can simplify this expansion yourself. We add one because when we subtract upper limit by lower limit, we exclude lower limit itself, so we have to add one to make it counts. See: how many term from 2 to 3? Two, right? just 2 and 3. Try different approach: \(3-2=1\) So we need to add one. Does that make sense?

OpenStudy (studygurl14):

I got the answer 60 too, but the expansion you gave above does not equal 60.... @geerky42

geerky42 (geerky42):

Check again.

OpenStudy (studygurl14):

Oops, sorry i hit the wrong button on my calculator. Yeah, it equals 60, lol.

OpenStudy (studygurl14):

Nice job @geerky42

OpenStudy (anonymous):

2,3,4,5,6 wasn't the right answer for the expansion. so , i dont get it . this is very frustrating. @geerky42

geerky42 (geerky42):

"2,3,4,5,6" is not right way to answer... I already gave you expansion. Now you just multiply number (don't add), then submit that.

OpenStudy (studygurl14):

The expansion was not 2,3,4,5,6.... @znae It was 3(2) + 3(3) + 3(4) + 3(5) + 3(6)

OpenStudy (anonymous):

oh ok, i see . thanks for the help y'all . @geerky42 @StudyGurl14

OpenStudy (studygurl14):

:)

OpenStudy (anonymous):

one more . \[\sum_{n=1}^{6}(n-1)\]

OpenStudy (studygurl14):

What are you doing? Solving, or expanding?

OpenStudy (anonymous):

expanding.

OpenStudy (studygurl14):

Do this: (1 - 1) + (2 -1) + (3 - 1) + (4 - 1) + (5 - 1) + (6 - 1) 0 + 1 + 2 + 3 + 4 + 5 @geerky42 Do you know whether or not the 0 is included?

OpenStudy (anonymous):

yes the 0 is included . thanks.